# Multi-Level Feature Aggregation and Recursive Alignment Network for Real-Time Semantic Segmentation

### v1

@@ Abstract
Real-time semantic segmentation is a crucial research for real-world applications. However, many methods lay particular emphasis on reducing the computational complexity and model size, while largely sacrificing the accuracy. In some scenarios, such as autonomous navigation and driver assistance system, accuracy and speed are equally important. To tackle this problem, we propose a novel Multi-level Feature Aggregation and Recursive Alignment Network (MFARANet), aiming to achieve high segmentation accuracy at real-time inference speed. We employ ResNet-18 as the backbone to ensure efficiency, and propose three core components to compensate for the reduced model capacity due to the shallow backbone. Specifically, we first design Multi-level Feature Aggregation Module (MFAM) to aggregate the hierarchical features in the encoder to each scale to benefit subsequent spatial alignment and multi-scale inference. Then, we build Recursive Alignment Module (RAM) by combining the flow-based alignment module with recursive upsampling architecture for accurate and efficient spatial alignment between multi-scale score maps. Finally, the Adaptive Scores Fusion Module (ASFM) is proposed to adaptively fuse multi-scale scores so that the final prediction can favor objects of multiple scales. Comprehensive experiments on three benchmark datasets including Cityscapes, CamVid and PASCAL-Context show the effectiveness and efficiency of our method. In particular, we achieve a better balance between speed and accuracy than state-of-the-art real-time methods on Cityscapes and CamVid datasets. Code is available at: https://github.com/Yanhua-Zhang/MFARANet
Index Terms:
@@
Real-time semantic segmentation, multi-scale inference, multi-level feature aggregation, spatial alignemnt.
I Introduction
@@
Semantic segmentation, aiming to label every pixel in the image, is one of the basic tasks of computer vision. Recently, architectural advances in Deep Convolutional Neural Networks (DCNNs) for image classification have greatly facilitated other visual recognition tasks[1, 2, 3], including semantic segmentation[4, 5, 6] and domain adaptation[7]. Although DCNNs-based semantic segmentation methods have significantly improved the accuracy, they are difficult to be directly applied to real-world applications due to large model size or high complexity. However, some applications, such as autonomous navigation or driver assistance systems, require segmentation models to maintain real-time inference speeds on large-size images to provide a wider field of view, placing a greater computational burden on the system. This makes real-time semantic segmentation a challenging research topic.
@@
To achieve real-time inference speed on high-end Graphics Processing Unit (GPU) cards, several models reduce computational complexity by adopting convolution factorization or designing shallow architectures. Among them, ERFNet[8] decomposes the standard 3 $\times$ 3 convolution into 3 $\times$ 1 and 1 $\times$ 3 operations by asymmetric convolutions, which theoretically reduces the computation by a factor of 1.5. ICNet [9] uses a shallow network to obtain spatial details, and inputs downsampled images into a deeper network for semantic information. Although above methods significantly improve inference speed, their accuracy is largely reduced compared to accuracy-oriented methods. For some real-world applications, the segmentation accuracy is also important.
@@
Obtaining multi-scale features is one of the key points to improve segmentation accuracy, as objects appear at various scales in the field of view [10, 11, 12]. Image pyramid is a straightforward way to obtain hierarchical features. It scales the original image to multiple resolutions to generate a set of images, which are then independently fed into the same neural network to obtain multi-scale scores [10, 13, 14]. Due to its effectiveness in promoting performance, image pyramid-based multi-scale inference has become a common operation in the evaluation phase of accuracy-oriented methods [15, 5]. However, the high computational complexity and large memory cost make it difficult to be directly used in real-world applications and end-to-end network training.
@@
Compared with image pyramid, utilizing the inherent multi-scale property of DCNNs is a more efficient way to obtain hierarchical features [2], which is widely used by accuracy-oriented [16, 17, 18] and speed-oriented segmentation networks [19, 20, 21]. By stacking strided convolution layers, different stages of the DCNNs extract multi-level features with different receptive fields and spatial resolutions. Previous works use this property in several different ways. One is to aggregate multi-stage features to the highest-resolution level, and then make predictions upon the fused feature map [22, 21, 23]. For example, DFA [20] designs a dual-path decoder to aggregate multi-level features of a lightweight backbone, while Panoptic FPN [24] directly fuses the upsampled features through element-wise sum operation. However, such feature-level fusion will lead to less efficiency of hierarchical representation due to semantic gaps and spatial misalignment.
@@
Instead of combining and predicting from feature pipeline, another way is to infer at each layer of the encoder to generate score maps, and then do a late fusion over them. The score maps (after segmentation heads, before the softmax function) are the log-probability of each category [10], so theoretically there is no semantic gap between multi-scale scores. Therefore, we believe that the score-level fusion can shrink the semantic gap existing in feature-level fusion. These methods either simply treat multi-level predictions as an additional precision boosting strategy [4, 12, 25], or focus on using the coarse to fine strategy [26, 27, 28]. By making predictions directly and independently at different stages of the encoder, they generate coarse or fine score maps based on the semantic or local spatial detail information contained in the corresponding features. It may be a suboptimal choice to obtain score maps at each level independently, because both high-level semantic information and low-level spatial details are important cues for segmentation [20, 21, 17]. Besides, when fusing score maps, the above methods simply upsample low-resolution maps by bilinear interpolation, which is difficult to be used to recover the spatial misalignment caused by long-distance connections and repeated downsampling [29, 30, 31].
@@
Differently, we design an architecture that provides rich hierarchical features for the inference at each level to obtain multi-scale predictions, each of which can better segment objects at a specific scale. In other words, by exploiting the inherent multi-scale property of DCNNs, our network performs multi-level inference and adaptive score fusion in a single-pass network to obtain the final prediction that favors various objects. To this end, we propose a Multi-level Feature Aggregation Module (MFAM) to aggregate hierarchical features of the encoder to each scale for further alignment and inference. Upon the features obtained from MFAM, we design a Recursive Alignment Module (RAM) to achieve efficient spatial alignment between multi-scale score maps. By combining the flow-based alignment module with our proposed recursive upsampling architecture, the RAM employs intermediate features for stepwise alignment. Furthermore, by adopting the pixel-wise attention to fuse multi-scale scores in an adaptive manner, the proposed Adaptive Scores Fusion Module (ASFM) generates the final prediction that can benefit the segmentation of objects of multiple scales. Finally, to enhance network training, we propose the Multi-scale Joint Supervision (MJS) to add extra supervision at each scale of the predictions. It employs Cross-Entropy (CE) to calculate the loss of segmentation prediction and boundary prediction, and uses a regularization term to maintain consistency between them.
@@
    To avoid spatial misalignment between scaled features, the RAM is designed to adopt intermediate features for stepwise alignment, which is more accurate and efficient than straightforward alignment. Furthermore, combined with MFAM, our RAM can align multi-scale score maps without being disturbed by semantic gaps.
@@
    Instead of fusing and inferring from the feature pipeline, we perform in-network score-level fusion to bridge the semantic gap existing in feature-level fusion, and also build the pixel-wise attention based ASFM to adaptively fuse multi-scale scores to improve accuracy.
@@
    Finally, by using the MJS to jointly supervise segmentation prediction and boundary prediction at each scale, we train our efficient network named MFARANet to perform multi-scale inference in a single-pass network. Taking full advantage of the inherent multi-scale properties of DCNNs, our architecture can compensate for the reduced model capacity due to the shallow backbone, enabling accurate real-time segmentation.
II Related Works
@@
Figure 1: Overall architecture of our approach. (a) The details of MFAM. ${S_{i}}$, ${D_{i}}$ and ${U_{i}}$ represent multi-level features from different stages of ResNet-18, Bottom-up path and Top-down path, respectively. The details of the “Feature Fusion” used in the two paths are shown in Fig. 2. (b) The process of aligning low-resolution features with the highest-resolution feature ${F_{1}}$. The detailed structure of RAM is illustrated in Fig. 4. ${F_{i}}$ indicates the i-th scale feature obtained from MFAM, and $P_{i}$ denotes the corresponding aligned feature. (c) The architecture of ASFM. “Seg Head” and “Attention Head” represent the modules for obtaining the score map (${\rm{Scor}}{{\rm{e}}_{i}}$) and weight map (${\rm{Weigh}}{{\rm{t}}_{i}}$) of the i-th scale, which are shown in detail in Fig. 5.
@@
There is a vast DCNNs-based literatures on accuracy-oriented or speed-oriented semantic segmentation. In the following, we mainly review the techniques of multi-scale inference, multi-level feature aggregation and feature alignment in semantic segmentation, which are related to our method.
II-A Multi-scale Inference
@@
Due to the high intra-class and inter-class scale variations of objects, obtaining multi-scale features or predictions is one of the key points to improve accuracy in semantic segmentation [12, 11, 2]. Performing multi-scale inference on the image pyramid is the most straightforward way to obtain multi-scale score maps. DeepLabv2 [32] extracts multiple score maps from a group rescaled original images through a shared network, and then performs cross scale fusion on them by maximum operations. Chen et al. [10] proposed using an attention mechanism to soft-weight each pixel between the score maps. Tao et al. [14] combined pixel-wise attention with the chain structure to fuse multi-scale predictions. At present, the multi-scale inference operation based on image pyramid has been widely used in the evaluation phase to boost segmentation accuracy [15, 5]. However, this method independently passes multiple scaled images through the network, which exponentially increases the computational complexity and memory consumption, making it difficult to be applied to end-to-end network training and real-time required applications.
@@
To promote efficiency, some networks perform inference on multi-level feature maps of the encoder to generate multiple score maps. FCN8s [4] and Hypercolumns [12] are early attempts to fuse the top-most prediction with middle-layer high-resolution predictions to generate the final prediction, with the aim of enabling their networks to learn finer details. Similarly, DAG-RNN [25] and DeepLab [33] directly employ this multi-level inference method as an additional performance boosting strategy. NDNet [34] uses convolutional layers to perform weighted fusion of multi-scale scores, and adapts separable convolutions to build a lightweight FCN8 to achieve real-time inference speed. GCN [26], CGBNet [27], and SABNet [28] all follow a similar network structure design pipeline and adopt a coarse to fine strategy, which utilizes high-resolution fine predictions from middle layers to gradually refine the topmost low-resolution coarse prediction. It may be a suboptimal choice to infer independently on single-level features of the encoder to obtain coarse or fine score maps, because multi-level features together provide rich cues for segmentation. Differently, we propose MFAM to aggregate multi-level features of the encoder to each scale, providing hierarchical information for further prediction. Besides, by using the RAM, we solve the spatial misalignment problem in the fusion process of multi-scale scores, which is ignored by the above methods.
II-B Multi-level Feature Aggregation
@@
As an effective method to improve accuracy, multi-level feature aggregation is used by both accuracy-oriented [24, 18] and speed-oriented segmentation methods [20, 35, 21]. The top-down path is used by FPN [2] to aggregate other levels of features to the lowest-level feature in an information flow manner, in which mid-level features are fused through lateral connections. Based on FPN-like architecture, lots of methods directly adopt element-wise sum [24] or channel concatenation [36, 18] to fuse multi-level features. Among the sophisticated methods, feature reuse is a commonly utilized mechanism. PANet [22] and SGCPNet [37] add a bottom-up path to integrate the middle layer features of the top-down path. In contrast, DFANet [20] and RGPNet [23] first employ the bottom-up path to aggregate features from the encoder, and then adopt a top-down path to refuse features of the bottom-up path. ShelfNet [35] uses three overlapping paths to reuse features from the previous path. For approaches different from feature reuse, Lin et al. [16] proposed an intertwining architecture to exchange information between adjacent scales in a bidirectional fashion. ZigZagNet [38] employs dense connections to blend feature maps between top-down and bottom-up networks. Weng et al. [21] used a long-distance skip connection to fuse the lowest-level feature with the middle-level features in the top-down path for real-time segmentation. All the above methods integrate multi-level features into the highest-resolution feature to obtain a single score map, while our MFAM aggregates multi-level features to each scale to generate multi-scale score maps.
II-C Feature Alignment/Upsampling

### v3

# Multi-Level Aggregation and Recursive Alignment Architecture for Efficient Parallel Inference Segmentation Network

@@ Abstract
Real-time semantic segmentation is a crucial research for real-world applications. However, many methods lay particular emphasis on reducing the computational complexity and model size, while largely sacrificing the accuracy. To tackle this problem, we propose a parallel inference network customized for semantic segmentation tasks to achieve a good trade-off between speed and accuracy. We employ a shallow backbone to ensure real-time speed, and propose three core components to compensate for the reduced model capacity to improve accuracy. Specifically, we first design a dual-pyramidal path architecture (Multi-level Feature Aggregation Module, MFAM) to aggregate multi-level features from the encoder to each scale, providing hierarchical clues for subsequent spatial alignment and corresponding in-network inference. Then, we build Recursive Alignment Module (RAM) by combining the flow-based alignment module with recursive upsampling architecture for accurate spatial alignment between multi-scale feature maps with half the computational complexity of the straightforward alignment method. Finally, we perform independent parallel inference on the aligned features to obtain multi-scale scores, and adaptively fuse them through an attention-based Adaptive Scores Fusion Module (ASFM) so that the final prediction can favor objects of multiple scales. Our framework shows a better balance between speed and accuracy than state-of-the-art real-time methods on Cityscapes and CamVid datasets. We also conducted systematic ablation studies to gain insight into our motivation and architectural design. Code is available at: https://github.com/Yanhua-Zhang/MFARANet.
keywords:
@@
Real-time semantic segmentation , parallel inference networks , multi-level feature aggregation , spatial alignment.
journal: arXiv
[…]
[1]organization=School of Astronautics, Northwestern Polytechnical University, addressline=127 West Youyi Road, city=Xi’an, postcode=710072, country=China
[…]
[2]organization=School of Artificial Intelligence, Optics and Electronics (iOPEN), Northwestern Polytechnical University, addressline=127 West Youyi Road, city=Xi’an, postcode=710072, country=China
[…]
[3]organization=Beijing Aerospace Automatic Control Research Institute, city=Beijing, postcode=100000, country=China
[…]
[4]organization=Xi’an University of Posts and Telecommunications, city=Xi’an, postcode=710072, country=China
1 Introduction
@@
Semantic segmentation, aiming to label every pixel in the image, is one of the basic tasks of computer vision. Recently, architectural advances in Deep Convolutional Neural Networks (DCNNs) for image classification have greatly facilitated other visual recognition tasks[1, 2, 3], including semantic segmentation[4, 5, 6]. Although DCNNs-based semantic segmentation methods have significantly improved the accuracy, they are difficult to be directly applied to real-world applications due to large model size or high complexity. Some applications, such as autonomous navigation or driver assistance systems, require segmentation models to maintain real-time inference speeds on large-size images to provide a wider field of view, placing a greater computational burden on the system. This makes real-time semantic segmentation a challenging research topic.
@@
To achieve real-time inference speed on high-end Graphics Processing Unit (GPU) cards, several models reduce computational complexity by adopting convolution factorization or designing shallow architectures. Among them, ERFNet[7] decomposes the standard 3 $\times$ 3 convolution into 3 $\times$ 1 and 1 $\times$ 3 operations by asymmetric convolutions, which theoretically reduces the computation by a factor of 1.5. ICNet [8] uses a shallow network to obtain spatial details, and inputs downsampled images into a deeper network for semantic information. Although above methods significantly improve inference speed, their accuracy is largely reduced. For some real-world applications, the segmentation accuracy is also important.
@@
Obtaining multi-scale features is one of the key points to improve segmentation accuracy, as objects appear at various scales in the field of view [9, 10, 11]. As shown in Fig. 2(a), image pyramid is a straightforward way to obtain hierarchical features. It scales the original image to multiple resolutions to generate a set of images, which are then independently fed into the same neural network to obtain multi-scale scores [9, 12, 13]. Due to its effectiveness in promoting performance, image pyramid-based multi-scale inference has become a common operation in the evaluation phase of accuracy-oriented methods [14, 15]. However, the high computational complexity and large memory cost make it difficult to be directly used in real-world applications and end-to-end network training.
@@
Compared with image pyramid, utilizing the inherent multi-scale property of DCNNs is a more efficient way to obtain hierarchical features [16], which is widely used by accuracy-oriented [17, 18, 19] and speed-oriented segmentation networks [20, 21, 22]. By stacking strided convolution layers, different stages of the DCNNs extract multi-level features with different receptive fields and spatial resolutions. In segmentation tasks, the most popular way to use this property is to aggregate multi-stage features to the highest-resolution level, and then make single-scale prediction upon the fused feature map [23, 22, 24]. For example, DFA [21] designs a dual-path decoder to aggregate multi-level features of a lightweight backbone, while Panoptic FPN [25] directly fuses the upsampled features through element-wise sum operation. However, such feature-level fusion will lead to less efficiency of hierarchical representation due to semantic gaps and spatial misalignment.
@@
Instead of combining and predicting from the feature pipeline, another way is to do parallel inference at each layer of the encoder to generate multiple score maps, and then do a late fusion over them. This in-network parallel inference architecture proposed by FPN [16] (Fig. 2(b)) is widely used in object detection tasks [26, 27, 28, 29]. However, the adaptation of this architecture to semantic segmentation tasks has not been deeply explored. Some segmentation networks either simply treat multi-level predictions as an additional precision boosting strategy [30, 11, 31], or focus on using the coarse to fine strategy [32, 33, 34]. They perform predictions directly and independently at different stages of the encoder, thus generating coarse or fine score maps based on the semantic or local spatial detail information contained in the corresponding features. Here, we point out that the FPN architecture designed for object detection is not suitable to be directly applied to segmentation tasks for two main reasons: firstly, it is a suboptimal choice to obtain score maps at each level independently, because both high-level semantic information and low-level spatial details are important cues for segmentation [21, 22, 18]. The image pyramid (Fig. 2(a)) performs inference of each scale on multi-level features, while FPN lacks lower-level features in the topmost and mid-scale inference; Secondly, FPN for object detection does not need to consider the fusion of multi-scale pixel-level predictions, which is an important factor for fine-grand segmentation tasks. When fusing score maps, the above mentioned segmentation methods simply upsample low-resolution maps by bilinear interpolation, which is difficult to be used to recover the spatial misalignment caused by long-distance connections and repeated downsampling [35, 36, 37].
Figure 1: Inference speed and Params vs. mIoU accuracy on the Cityscapes test set. The size of the points indicates the model size. Purple points are representative real-time methods in Table. 9. The red and green points represent our MFARANet using 1024 $\times$ 1024 or the whole image (1024 $\times$ 2048) input for inference, respectively. ’Pruned’ indicates the special network pruning method used for our parallel inference network, which is shown in Fig. 9 and discussed in Section. 4.2.8. All FPS are measured on a single GTX 3090 GPU with the image resolution on which the inference is performed to calculate the accuracy.
(a) Image Pyramid
(b) Feature Pyramid Networks
(c) Our Network
Figure 2: The comparison between different parallel inference networks. The boxes represent feature maps, and their length and number roughly reflect the relative spatial resolution between them. (a) Independently passing the scaled images through the whole segmentation network (omitted for simplicity) to obtain multi-scale predictions. (b) The in-network parallel inference architecture proposed by Feature Pyramid Networks (FPN) [16] for object detection tasks. (c) Our custom-designed parallel inference network for fine-grained real-time semantic segmentation tasks, which is much faster than image pyramid.
@@
Differently, we design a parallel inference network customized for semantic segmentation tasks that can provide rich hierarchical features for the prediction at each level to obtain multi-scale scores as if obtained from image pyramid while being more efficient than it. To this end, we propose a Multi-level Feature Aggregation Module (MFAM) to aggregate hierarchical features of the encoder to each scale for further alignment and inference. Upon the features obtained from MFAM, we design a Recursive Alignment Module (RAM) to achieve spatial alignment between multi-scale score maps, which has half the computational complexity of the straightforward alignment method. By combining the flow-based alignment module with our proposed recursive upsampling architecture, the RAM employs intermediate features for stepwise alignment. Furthermore, by adopting the pixel-wise attention to fuse multi-scale scores in an adaptive manner, the proposed Adaptive Scores Fusion Module (ASFM) generates the final prediction that can benefit the segmentation of objects of multiple scales. It is worth mentioning that from the ablation experiments, we found a pruning method specific to this parallel inference architecture, that is, training on the entire network and removing certain scales during testing phrase. However, using network pruning to reduce model complexity is not the focus of this work, so it is only illustrated in Fig. 9 and discussed in Section. 4.2.8. Finally, to enhance network training, we propose the Multi-scale Joint Supervision (MJS) to add extra supervision at each scale of the predictions. It employs Cross-Entropy (CE) to calculate the loss of segmentation prediction and boundary prediction, and uses a regularization term to maintain consistency between them.
@@
    To avoid spatial misalignment between scaled features, the RAM is designed to adopt intermediate features for stepwise alignment, which is more accurate and faster than straightforward alignment. Furthermore, combined with MFAM, our RAM can align multi-scale score maps without being disturbed by semantic gaps.
@@
    Instead of fusing and inferring from the feature pipeline, we design a parallel inference network customized for semantic segmentation tasks to obtain multi-scale scores, and also build the pixel-wise attention based ASFM to adaptively fuse them to improve accuracy.
@@
    Finally, we use the MJS to jointly supervise segmentation prediction and boundary prediction at each scale to enhance feature representation and boost network training.
2 Related Works
@@
Figure 3: Overall architecture of our approach. (a) The details of MFAM. ${S_{i}}$, ${D_{i}}$ and ${U_{i}}$ represent multi-level features from different stages of ResNet-18, Bottom-up path and Top-down path, respectively. (b) The process of aligning low-resolution features with the highest-resolution feature ${F_{1}}$. The detailed structure of RAM is illustrated in Fig. 4. ${F_{i}}$ indicates the i-th scale feature obtained from MFAM, and $P_{i}$ denotes the corresponding aligned feature. (c) The architecture of ASFM. “Seg Head” and “Attention Head” represent the modules for obtaining the score map (${\rm{Scor}}{{\rm{e}}_{i}}$) and weight map (${\rm{Weigh}}{{\rm{t}}_{i}}$) of the i-th scale, which are shown in detail in Fig. 6.
@@
There is a vast DCNNs-based literatures on accuracy-oriented or speed-oriented semantic segmentation. In the following, we mainly review the techniques of multi-scale/parallel inference, multi-level feature aggregation and feature alignment in semantic segmentation, which are related to our method.
2.1 Multi-scale/Parallel Inference
@@
Due to the high intra-class and inter-class scale variations of objects, obtaining multi-scale features or predictions is one of the key points to improve accuracy in semantic segmentation [11, 10, 16]. Performing multi-scale inference on the image pyramid (Fig. 2(a)) is the most straightforward way to obtain multi-scale score maps. DeepLabv2 [38] extracts multiple score maps from a group rescaled original images through a shared network, and then performs cross scale fusion on them by maximum operations. Chen et al. [9] proposed using an attention mechanism to soft-weight each pixel between the score maps. Tao et al. [13] combined pixel-wise attention with the chain structure to fuse multi-scale predictions. At present, the multi-scale inference operation based on image pyramid has been widely used in the evaluation phase to boost segmentation accuracy [14, 15]. However, this method independently passes multiple scaled images through the network, which exponentially increases the computational complexity and memory consumption, making it difficult to be applied to end-to-end network training and real-time required applications.
@@
To promote efficiency, some networks perform parallel inference on multi-level feature maps of the encoder to generate multiple score maps. FCN8s [30] and Hypercolumns [11] are early attempts to fuse the top-most prediction with middle-layer high-resolution predictions to generate the final prediction, with the aim of enabling their networks to learn finer details. Similarly, DAG-RNN [31] and DeepLab [39] directly employ this multi-level parallel inference method as an additional performance boosting strategy. NDNet [40] uses convolutional layers to perform weighted fusion of multi-scale scores, and adapts separable convolutions to build a lightweight FCN8 to achieve real-time inference speed. GCN [32], CGBNet [33], and SABNet [34] all follow a similar network structure design pipeline and adopt a coarse to fine strategy, which utilizes high-resolution fine predictions from middle layers to gradually refine the topmost low-resolution coarse prediction.
It is a suboptimal choice to infer independently on single-level features of the encoder to obtain coarse or fine score maps, because multi-level features together provide rich cues for segmentation. Differently, we propose MFAM to aggregate multi-level features of the encoder to each scale, providing hierarchical information for further prediction. Besides, by using the RAM, we solve the spatial misalignment problem in the fusion process of multi-scale scores, which is ignored by the above methods.
2.2 Multi-level Feature Aggregation
@@
As an effective method to improve accuracy, multi-level feature aggregation is used by both accuracy-oriented [25, 19] and speed-oriented segmentation methods [21, 41, 22]. The top-down path is used by FPN [16] (Fig. 2(b)) to aggregate other levels of features to the lowest-level feature in an information flow manner, in which mid-level features are fused through lateral connections. Based on FPN-like architecture, lots of methods directly adopt element-wise sum [25] or channel concatenation [42, 19] to fuse multi-level features. Among the sophisticated methods, feature reuse is a commonly utilized mechanism. PANet [23] and SGCPNet [43] add a bottom-up path to integrate the middle layer features of the top-down path. In contrast, DFANet [21] and RGPNet [24] first employ the bottom-up path to aggregate features from the encoder, and then adopt a top-down path to refuse features of the bottom-up path. ShelfNet [41] uses three overlapping paths to reuse features from the previous path. For approaches different from feature reuse, Lin et al. [17] proposed an intertwining architecture to exchange information between adjacent scales in a bidirectional fashion. ZigZagNet [44] employs dense connections to blend feature maps between top-down and bottom-up networks. Weng et al. [22] used a long-distance skip connection to fuse the lowest-level feature with the middle-level features in the top-down path for real-time segmentation. All the above methods integrate multi-level features into the highest-resolution feature to obtain a single score map, while our MFAM aggregates multi-level features to each scale to generate multi-scale score maps.
2.3 Feature Alignment/Upsampling